# Chern character of Schur bundles

### v1

@@ Abstract.
Given a vector bundle $E$, we give an explicit formula to compute Chern classes of Schur bundles $\mathsf{\mathbf{S}}^{\alpha}E$ in terms of those of $E$.
@@ Introduction
where $\alpha$ is a Young diagram of length at most $\operatorname{rk}(E)$. If we denote by $V$ a fibre of $E$, then a fibre of $\mathsf{\mathbf{S}}^{\alpha}E$ is the irreducible Schur $\operatorname{GL}(V)$-module $\mathsf{\mathbf{S}}^{\alpha}V$.
An important problem in computational algebraic geometry is to compute the Chern classes of $\mathsf{\mathbf{S}}^{\alpha}E$ in terms of the Chern classes of $E$. It is in principle possible to do such a computation by using the splitting principle, but the combinatorics behind this approach is very cumbersome, see for example [Svr93], where the case of symmetric and alternating powers was treated. In this short paper we provide a simple solution to this problem.
Theorem (Theorem 3.2).
For any $n\geq 0$, the $n^{\operatorname{th}}$-Chern character coefficient of $\mathsf{\mathbf{S}}^{\alpha}E$ is
@@
  • $\operatorname{ch}^{(\mu)}(E)=\frac{\operatorname{ch}_{1}(E)^{\mu_{1}}}{\mu_{1}!}\frac{\operatorname{ch}_{2}(E)^{\mu_{2}}}{\mu_{2}!}\cdots$ are the divided powers of the Chern character coefficients.
The quantity $\mathsf{D}^{\nu}(\mathsf{\mathbf{s}}_{\alpha})$, the new input in the formula, is a purely combinatorial object which is easily computed in explicit examples, see Examples 3.5 and 3.6 and Proposition 3.8.
The proof of Theorem 3.2 is simple and only uses some weak properties of the Chern character, like multiplicativity and additivity. To emphasise this aspect, and thus making the proof more transparent, we undress it of all the superficial structures and reduce it to a statement about polynomials in countably many variables. This is the content of our Key Lemma (Lemma 1.3), from which all the results in the paper are deduced.
Notice that this abstract point of view is not only an exercise of style. The same proof in fact ensures that the result holds in any category where suitable characteristic classes can be defined. For example, it holds for proper algebraic varieties and locally free sheaves over them. Similarly, it also holds for characteristic classes valued in a Chow ring, instead of a cohomology ring. As an example of characteristic classes that differs from the Chern character, we also prove the analog statement for the twisted Chern character $k(F)=\exp\left(-\frac{c_{1}(F)}{\operatorname{rk}(F)}\right)\operatorname{ch}(F)$ considered in [Mar20] as characteristic classes of projective bundles (see Theorem 4.1).
As already mentioned, for symmetric and exterior powers a concrete formula was found by Svrtan in [Svr93]. We give alternative formulas for these special Schur bundles in Proposition 3.8 (see Remark 3.11). The formula for the first Chern class of a Schur bundle was a folklore result, whose first rigorous proof can be found in [Rub13]. We present in Corollary 3.3 an alternative proof.
In our previous work [DFO25a] (see also [DFO25b, Appendix A]) we computed the second and third Chern characters of Schur functors in a similar spirit to the one presented here. This work extends and complements those results, see § 4.
@@
The technical core of the paper is contained in § 1, where the Key Lemma is proved. In § 2 we recover some important facts about the ring of symmetric functions, that will be used in § 3, where the Theorem is proved. In § 4 we make explicit the example of the twisted Chern character.
@@ Acknowledgements
This research has been partially funded by the European Union - NextGenerationEU under the National Recovery and Resilience Plan (PNRR) - Mission 4 Education and research - Component 2 From research to business - Investment 1.1 Notice Prin 2022 - DD N. 104 del 2/2/2022, from title “Symplectic varieties: their interplay with Fano manifolds and derived categories”, proposal code 2022PEKYBJ – CUP J53D23003840006. The second and third authors are members of the INDAM-GNSAGA group.
@@ 1. The Key Lemma
Let $R=\mathbb{Q}[x_{1},x_{2},\dots]$ be a polynomial ring in countably many variables $x_{k}$, $k\geq 1$; we think of $R=\bigoplus_{d\geq 0}R_{d}$ as a graded ring by setting the degree of each indeterminate to be $1$. We are also given a commutative unital graded $\mathbb{Q}$-algebra $A=\bigoplus_{n\geq 0}A_{n}$, where $\mathbb{Q}\subset A_{0}$, whose completion we denote by $\widehat{A}=\prod_{n\geq 0}A_{n}$. Finally, we fix a collection of positive rational numbers $\mathsf{q}=\{q_{n,k}\}_{n\geq 1,\,k\geq 1}$ with $q_{n,1}=1$.
@@
Consider a map $\mathsf{C}:R\to\widehat{A}$ whose projection to each graded component we denote by $\mathsf{C}_{n}\colon R\to A_{n}$, so that
@@
    $\mathsf{C}$ is multiplicative, i.e. $\mathsf{C}(fg)=\mathsf{C}(f)\mathsf{C}(g)$ for all $f,g\in\mathbb{Q}[x_{1},x_{2},\dots]$;
@@
We refer to the property in (2) above as same-degree additivity. For our main application, $\mathsf{C}$ will be the Chern character, which is additive; nevertheless, allowing this more general setting gives us the possibility to extend the results to other classes, see for example § 4.
For each $n\geq 0$, let us now consider the derivation
$\mathsf{D}_{n}=\sum_{k\geq 1}q_{n,k}\frac{\partial}{\partial x_{k}}\in\operatorname{Der}R.$ (1.1)
Notice that $\mathsf{D}_{n}$ is locally nilpotent and well-defined, since for each polynomial $f\in\mathbb{Q}[x_{1},x_{2},\dots]$ only finitely many of the $\frac{\partial f}{\partial x_{k}}$ are non-zero.
The formal sum $\sum_{n\geq 1}\mathsf{C}_{n}(x_{1})\cdot\mathsf{D}_{n}$ is then a (locally nilpotent, converging) derivation of $R_{\widehat{A}}=\widehat{A}\otimes_{\mathbb{Q}}R=\widehat{A}[x_{1},x_{2},\dots]$ and its exponential is a well defined automorphism of $R_{\widehat{A}}$, whose restriction to $R\subset R_{\widehat{A}}$ yields a $\mathbb{Q}$-algebra homomorphism
$\exp\left(\sum_{n\geq 1}\mathsf{C}_{n}(x_{1})\cdot\mathsf{D}_{n}\right)\colon\mathbb{Q}[x_{1},x_{2},\dots]\longrightarrow R_{\widehat{A}}=\widehat{A}\otimes_{\mathbb{Q}}\mathbb{Q}[x_{1},x_{2},\dots].$
@@
$\mathsf{C}_{0}\circ\exp\left(\sum_{n\geq 1}\mathsf{C}_{n}(x_{1})\cdot\mathsf{D}_{n}\right)\colon R\longrightarrow\widehat{A}$
@@
Let $\mathsf{C}$ be a $\widehat{A}-$valued $\mathsf{q}$-characteristic class. Then,
$\mathsf{C}=\mathsf{C}_{0}\circ\exp\left(\sum_{n\geq 1}\mathsf{C}_{n}(x_{1})\cdot\mathsf{D}_{n}\right),$ (1.2)
when both sides are applied to any given homogeneous polynomial in $R$. If furthermore $\mathsf{C}$ is a $\mathbb{Q}$-algebra homomorphism, then equality (1.2) holds in general.
@@
As for the right-hand side, we notice that $\mathsf{D}_{n}(x_{k})=q_{n,k}\in\mathbb{Q}$, so that $\mathsf{D}_{m}(\mathsf{D}_{n}(x_{k}))=0$ for all $m$ and $n$. When we apply the automorphism to $x_{k}$ only the constant and the linear term survive and we obtain
$\exp\left(\sum_{n\geq 1}\mathsf{C}_{n}(x_{1})\cdot\mathsf{D}_{n}\right)(x_{k})=x_{k}+\sum_{n\geq 1}q_{n,k}\mathsf{C}_{n}(x_{1}).$
@@
By multiplicativity of $\mathsf{C}$, Equation (1.2) holds for each monomial, and since the restriction of $\mathsf{C}$ on each graded component of $R$ is $\mathbb{Q}$-linear, it also holds for every homogeneous polynomial. ∎
For the following corollary we use the notation $\nu=(1^{\nu_{1}},2^{\nu_{2}},\dots,h^{\nu_{h}})$ for partitions. In particular, $\nu\vdash n$ means that $\sum_{i}i\nu_{i}=n$.
If $\nu=(1^{\nu_{1}},2^{\nu_{2}},\dots,h^{\nu_{h}})$ is a partition, then we put
$\mathsf{D}^{\nu}=D_{1}^{\nu_{1}}\circ\cdots\circ\mathsf{D}_{h}^{\nu_{h}}.$ (1.3)
Notice that, for any $n,m\in\mathbb{Z}$, the operators $\mathsf{D}_{n}$ and $\mathsf{D}_{m}$ commute, i.e. $\mathsf{D}_{n}\mathsf{D}_{m}=\mathsf{D}_{m}\mathsf{D}_{n}$.
@@
Let $\mathsf{C}$ be a $\hat{A}-$valued $\mathsf{q}$-characteristic class. Then for every $n\geq 0$ and every $f\in\mathbb{Q}[x_{1},x_{2},\dots]$
$\mathsf{C}_{n}(f)=\sum_{\nu\vdash n}\mathsf{C}_{0}(\mathsf{D}^{\nu}(f))\mathsf{C}^{(\nu)}(x_{1}),$
where $\mathsf{C}^{(\nu)}=\prod_{i}\frac{\mathsf{C}_{i}^{\nu_{i}}}{\nu_{i}!}$ is the divided power of $\mathsf{C}$. ∎
@@ 2. Symmetric functions
In this short section we recall some facts about symmetric functions, see [Mac95] for a reference.
@@
  • Given $(i,j)\in\alpha$, we denote by $\operatorname{Hook}_{(i,j)}(\alpha)$ the hook centred in position $(i,j)$ and by $\operatorname{h}_{(i,j)}(\alpha)$ he corresponding hook number, i.e. its size.
  • Partitions will be denoted by $\mu=(1^{\mu_{1}},2^{\mu_{2}},\dots)$ and their size is $|\mu|=\sum_{i}i\mu_{i}$.
  • $\Lambda=\bigoplus_{d}\Lambda^{d}$ is the graded ring of symmetric functions with coefficients in $\mathbb{Q}$.
@@
  • Power sums functions are $\mathsf{\mathbf{p}}_{k}(x_{1},x_{2},\dots)=\sum_{i\geq 1}x_{i}^{k}\in\Lambda^{k}$.
  • If $\mu$ is a partition, then we set
    $\mathsf{\mathbf{p}}_{\mu}:=\mathsf{\mathbf{p}}_{1}^{\mu_{1}}\mathsf{\mathbf{p}}_{e}^{\mu_{2}}\cdots\in\Lambda^{|\mu|}$
    for the corresponding monomial in power sums.
@@
Let $\Lambda=\bigoplus\Lambda^{d}$ be the graded ring of symmetric functions with coefficients in $\mathbb{Q}$. Two sets of generators of $\Lambda$ are given by Schur functions $\mathsf{\mathbf{s}}_{\alpha}$ and power sums functions $\mathsf{\mathbf{p}}_{k}$. A Schur function $\mathsf{\mathbf{s}}_{\alpha}$ is homogeneous of degree $|\alpha|$, while a power sums function $\mathsf{\mathbf{p}}_{k}$ is homogeneous of degree $k$.
The homogeneous component $\Lambda^{d}$ is generated (as a $\mathbb{Q}$-vector space) by Schur functions $\mathsf{\mathbf{s}}_{\alpha}$ with $|\alpha|=d$ or by monomials in power sums $\mathsf{\mathbf{p}}_{\mu}$ with $|\mu|=d$.
There is a natural isomorphism
$\Lambda\cong\mathbb{Q}[\mathsf{\mathbf{p}}_{1},\mathsf{\mathbf{p}}_{2},\dots]$
exhibiting the ring of symmetric functions as a polynomial ring in countably many variables.
@@
For any integer $m$, define the derivations
$\mathsf{D}_{m}\colon\Lambda\to\Lambda\,,\qquad\mathsf{D}_{m}=\sum_{k\geq 0}k^{m}\frac{\partial}{\partial\mathsf{\mathbf{p}}_{k}}.$ (2.1)
In [DFO25b] we considered a derivation $\mathsf{D}_{s}$, where $s$ is a formal variable. There it was showed to be a quasi-isometry, property that allowed us to give recursive formulas to compute tensor products and plethysms of Schur functions. Here we are considering the evaluation of $\mathsf{D}_{s}$ at $s=m$.
Next, we want to explain how to explicitly compute $\mathsf{D}_{m}(\mathsf{\mathbf{s}}_{\alpha})$. The following interpretation is proved in [DFO25b, Section 3] (see also [DFO25b, Proposition 2.10]).
For each $(i,j)\in\alpha$ we denote by $\alpha_{(i,j)}$ the Young diagram obtained by $\alpha$ by removing the hook $\operatorname{Hook}_{(i,j)}(\alpha)$. If removing a hook disconnects the diagram, then we move the lower component up-left.
@@
$\mathsf{D}_{m}(\mathsf{\mathbf{s}}_{\alpha})=\sum_{(i,j)\in\alpha}(-1)^{\varepsilon(i,j)}\mathsf{\mathbf{s}}_{\alpha_{(i,j)}}\operatorname{h}_{(i,j)}^{m-1},$
@@
$\mathsf{D}_{m}(\mathsf{\mathbf{s}}_{(3,2,2)})=\mathsf{\mathbf{s}}_{(3,2,1)}+\mathsf{\mathbf{s}}_{2,2,2}+(\mathsf{\mathbf{s}}_{(3,2)}-\mathsf{\mathbf{s}}_{(3,1,1)})2^{m-1}-\mathsf{\mathbf{s}}_{3,1}3^{m-1}+\mathsf{\mathbf{s}}_{1,1,1}4^{m-1}+\mathsf{\mathbf{s}}_{(1,1)}5^{m-1}\,.$
@@
$\mathsf{D}_{m}(\mathsf{\mathbf{s}}_{n})=\sum_{k=1}^{n}k^{m-1}\mathsf{\mathbf{s}}_{n-k}\qquad\mbox{and}\qquad\mathsf{D}_{m}(\mathsf{\mathbf{s}}_{1^{n}})=\sum_{k=1}^{n}(-1)^{k+1}k^{m-1}\mathsf{\mathbf{s}}_{1^{n-k}}.$
@@
If $|\alpha|=n$, then $\mathsf{D}^{\mu}(\mathsf{\mathbf{s}}_{\alpha})=0$ for every $\mu=(1^{\mu_{1}},\dots,\ell^{\mu_{\ell}})$ such that $\mu_{1}+\cdots+\mu_{l}>n$. (If we express $\mu$ as a Young diagram, then $\mu_{1}+\cdots+\mu_{l}$ is the number of rows of the diagram.)
3. Chern character of Schur bundles
Through this section we assume that $X$ is a connected compact complex manifold. In this case we consider, as graded commutative algebra $A$, the even cohomology ring $\operatorname{H}^{2\ast}(X,\mathbb{Q})$. Notice that since $\operatorname{H}^{2n}(X,\mathbb{Q})\neq 0$ only for finitely many $n$, we have $A=\widehat{A}$.

### v2

@@ Abstract.
We provide explicit formulas for computing the Chern character of Schur bundles $\mathsf{\mathbf{S}}^{\alpha}E$ in terms of that of $E$.
2020 Mathematics Subject Classification:
Primary 19L10, 05E10; Secondary 11B68, 20G05
@@ Introduction
where $\alpha$ is a Young diagram of length at most $\operatorname{rk}(E)$. If we denote by $V$ any fibre of $E$, then the corresponding fibre of $\mathsf{\mathbf{S}}^{\alpha}E$ is the irreducible Schur $\operatorname{GL}(V)$-module $\mathsf{\mathbf{S}}^{\alpha}V$.
An important problem in computational algebraic geometry is to compute Chern classes of $\mathsf{\mathbf{S}}^{\alpha}E$ in terms of those of $E$. It is in principle possible to perform such computations using the splitting principle, but the combinatorics behind this approach is very cumbersome, see for example [Svr93, Man94], where the cases of symmetric and alternating powers were treated. In this short paper we provide a compact and effective solution to this problem.
Theorem A (Theorem 3.4).
For any $n\geq 0$, the $n^{\operatorname{th}}$ Chern character coefficient of $\mathsf{\mathbf{S}}^{\alpha}E$ is
@@
  • $\operatorname{ch}^{(\nu)}(E)=\frac{\operatorname{ch}_{1}(E)^{\nu_{1}}}{\nu_{1}!}\frac{\operatorname{ch}_{2}(E)^{\nu_{2}}}{\nu_{2}!}\cdots$ are the divided powers of the Chern character coefficients.
The quantity $\mathsf{D}^{\nu}(\mathsf{\mathbf{s}}_{\alpha})$, which is the new ingredient in the formula, is a purely combinatorial object which is easily computed in explicit examples, see Examples 3.6 and 3.7 and Proposition 4.2.
The proof of Theorem 3.4 is simple and only uses standard properties of the Chern character, such as multiplicativity and additivity. To emphasise this aspect, thus making the proof more transparent, we undress it of all the superficial structures and reduce it to a statement about polynomials in countably many variables. This is the content of our key Lemma 1.3, from which all the results in the paper are deduced.
Notice that this abstract point of view is not only an exercice de style. The same proof in fact ensures that the result holds in any category where suitable characteristic classes can be defined. For example, it holds for proper algebraic varieties and locally free sheaves over them. Similarly, it also holds for characteristic classes valued in a Chow ring, instead of a cohomology ring. As an example of characteristic classes that differs from the Chern character, we also prove the analogous statement for the twisted Chern character
$k(F)=\exp\left(-\frac{c_{1}(F)}{\operatorname{rk}(F)}\right)\operatorname{ch}(F)$
considered in [Mar20] as characteristic classes of projective bundles.
Theorem B (Theorem 3.9).
For any $n\geq 0$, the $n^{\operatorname{th}}$ twisted Chern character coefficient of $\mathsf{\mathbf{S}}^{\alpha}E$ is
$k_{n}(\mathsf{\mathbf{S}}^{\alpha}E)=\sum_{\nu\vdash n}\operatorname{rk}(\mathsf{D}^{\nu}(\mathsf{\mathbf{s}}_{\alpha}))\,k^{(\nu)}(E),$
where the sum may be taken over the partitions of $n$ into parts that are all strictly larger than $1$.
In our previous work [DFO25a] we proved formulas expressing $k_{2}(\mathsf{\mathbf{S}}^{\alpha}E)$ and $k_{3}(\mathsf{\mathbf{S}}^{\alpha}E)$ as certain multiples of $k_{2}(E),k_{3}(E)$. Theorem B reinterprets those results as
$k_{2}(\mathsf{\mathbf{S}}^{\alpha}E)=$ $\,\operatorname{rk}(\mathsf{D}_{2}(\mathsf{\mathbf{s}}_{\alpha}))\,k_{2}(E)\,$,
$k_{3}(\mathsf{\mathbf{S}}^{\alpha}E)=$ $\,\operatorname{rk}(\mathsf{D}_{3}(\mathsf{\mathbf{s}}_{\alpha}))\,k_{3}(E)\,$,
and provides generalizations to all twisted Chern character coefficients $k_{n}$, see § 3.2.
As already mentioned, concrete formulas for the Chern character of symmetric and exterior powers were found by Svrtan in [Svr93] and Manivel in [Man94, Section 2]. We recover those formulas in Proposition 4.2 and Proposition 4.3, and extend them to twisted Chern characters.
The formula for the first Chern class of a Schur bundle has long been folklore knowledge, and its first rigorous proof can be found in [Rub13]. We present in Corollary 3.5 an alternative short proof.
@@
The technical core of the paper is contained in § 1, where the Key Lemma is proved. In § 2 we recover some important facts about the ring of symmetric functions, that will be used in § 3, where Theorems A and B are proved. Finally, in § 4 we prove formulas for Chern character and twisted Chern character of symmetric and exterior powers.
@@ Acknowledgements
We are glad to thank Laurent Manivel for comments on a first draft of the paper. This research has been partially funded by the European Union - NextGenerationEU under the National Recovery and Resilience Plan (PNRR) - Mission 4 Education and research - Component 2 From research to business - Investment 1.1 Notice Prin 2022 - DD N. 104 del 2/2/2022, from title “Symplectic varieties: their interplay with Fano manifolds and derived categories”, proposal code 2022PEKYBJ – CUP J53D23003840006. The second and third authors are members of the INDAM-GNSAGA group.
@@ 1. The Key Lemma
Let $R=\mathbb{Q}[x_{1},x_{2},\dots]$ be a polynomial ring with an explicit choice of countably many free commutative $\mathbb{Q}$-algebra generators $x_{k}$, $k\geq 1$; we think of $R=\bigoplus R_{d}$ as a graded ring in which each indeterminate $x_{i}$ is homogeneous of any arbitrarily chosen degree. [^1] The only role played by the grading of $R$ is in extending our statements below from monomials to all homogeneous polynomials, so the nature of the grading or the monoid/group used to grade $R$ are not really relevant. We are also given a commutative unital $\mathbb{N}$-graded $\mathbb{Q}$-algebra $A=\bigoplus_{n\geq 0}A_{n}$, where $\mathbb{Q}\subset A_{0}$, whose completion we denote by $\widehat{A}=\prod_{n\geq 0}A_{n}$. Finally, we fix a collection of positive rational numbers $\mathsf{q}=\{q_{n,k}\}_{n\geq 1,\,k\geq 1}$ with $q_{n,1}=1$.
@@
Consider a map $\mathsf{C}:R=\mathbb{Q}[x_{1},x_{2},\dots]\to\widehat{A}$ whose projection to each graded component we denote by $\mathsf{C}_{n}\colon R\to A_{n}$, so that
@@
    $\mathsf{C}$ is multiplicative, i.e. $\mathsf{C}(fg)=\mathsf{C}(f)\mathsf{C}(g)$ for all $f,g\in R$;
@@
We refer to property (3) above as same-degree additivity. In our main application $\mathsf{C}$ will be the Chern character, which is additive; nevertheless, allowing this more general setting gives us the possibility to extend the results to other classes, see for example § 3.2.
For each $n\geq 1$, let us now consider the derivations
$\delta_{\mathsf{q},n}=\sum_{k\geq 1}q_{n,k}\frac{\partial}{\partial x_{k}}\in\operatorname{Der}R.$ (1.1)
Notice that each $\delta_{\mathsf{q},n}$ is locally nilpotent and well defined, since for each polynomial $f\in\mathbb{Q}[x_{1},x_{2},\dots]$ only finitely many of the $\frac{\partial f}{\partial x_{k}}$ are non-zero.
The formal sum $\sum_{n\geq 1}\mathsf{C}_{n}(x_{1})\cdot\delta_{\mathsf{q},n}$ is then a (locally nilpotent, converging) derivation of $R_{\widehat{A}}=\widehat{A}\otimes_{\mathbb{Q}}R=\widehat{A}[x_{1},x_{2},\dots]$ and its exponential is a well defined automorphism of $R_{\widehat{A}}$, whose restriction to $R\subset R_{\widehat{A}}$ yields a $\mathbb{Q}$-algebra homomorphism
$\exp\left(\sum_{n\geq 1}\mathsf{C}_{n}(x_{1})\cdot\delta_{\mathsf{q},n}\right)\colon\mathbb{Q}[x_{1},x_{2},\dots]\longrightarrow R_{\widehat{A}}=\widehat{A}\otimes_{\mathbb{Q}}\mathbb{Q}[x_{1},x_{2},\dots].$
@@
$\mathsf{C}_{0}\circ\exp\left(\sum_{n\geq 1}\mathsf{C}_{n}(x_{1})\cdot\delta_{\mathsf{q},n}\right)\colon R\longrightarrow\widehat{A}$
@@
Let $\mathsf{C}$ be an $\widehat{A}-$valued $\mathsf{q}$-characteristic class. Then equality
$\mathsf{C}=\mathsf{C}_{0}\circ\exp\left(\sum_{n\geq 1}\mathsf{C}_{n}(x_{1})\cdot\delta_{\mathsf{q},n}\right)$ (1.2)
holds whenever both sides are applied to any given homogeneous polynomial in $R$. If furthermore $\mathsf{C}$ is a $\mathbb{Q}$-algebra homomorphism, then equality (1.2) holds in general.
@@
As for the right-hand side, we notice that $\delta_{\mathsf{q},n}(x_{k})=q_{n,k}\in\mathbb{Q}$, so that $\delta_{\mathsf{q},m}(\delta_{\mathsf{q},n}(x_{k}))=0$ for all $m$ and $n$. When we apply the automorphism to $x_{k}$, only the constant and the linear term survive and we obtain
$\exp\left(\sum_{n\geq 1}\mathsf{C}_{n}(x_{1})\cdot\delta_{\mathsf{q},n}\right)(x_{k})=x_{k}+\sum_{n\geq 1}q_{n,k}\mathsf{C}_{n}(x_{1}).$
@@
By multiplicativity of $\mathsf{C}$, Equation (1.2) holds for each monomial, and since the restriction of $\mathsf{C}$ on each graded component of $R$ is $\mathbb{Q}$-linear, it also holds for every homogeneous polynomial. The final claim is clear. ∎
For the following corollary we use the notation $\nu=(1^{\nu_{1}},2^{\nu_{2}},\dots,h^{\nu_{h}})$ for partitions. In particular, $\nu\vdash n$ means that $\sum_{i}i\nu_{i}=n$. Then we set
$\delta_{\mathsf{q}}^{\nu}=\delta_{\mathsf{q},1}^{\nu_{1}}\circ\cdots\circ\delta_{\mathsf{q},h}^{\nu_{h}}.$ (1.3)
Notice that, for any $n,m\in\mathbb{Z}$, the operators $\delta_{\mathsf{q},n}$ and $\delta_{\mathsf{q},m}$ commute, i.e. $\delta_{\mathsf{q},n}\delta_{\mathsf{q},m}=\delta_{\mathsf{q},m}\delta_{\mathsf{q},n}$.
@@
Let $\mathsf{C}$ be a $\hat{A}-$valued $\mathsf{q}$-characteristic class. Then for every $n\geq 0$ and every homogeneous $f\in\mathbb{Q}[x_{1},x_{2},\dots]$
$\mathsf{C}_{n}(f)=\sum_{\nu\vdash n}\mathsf{C}_{0}(\delta_{\mathsf{q}}^{\nu}(f))\mathsf{C}^{(\nu)}(x_{1}),$
where $\mathsf{C}^{(\nu)}=\prod_{i}\frac{\mathsf{C}_{i}^{\nu_{i}}}{\nu_{i}!}$ denotes divided powers of coefficients of $\mathsf{C}$. ∎
@@ 2. Symmetric functions
In this short section we recall standard facts about symmetric functions, see [Mac95] for a reference.
@@
  • Given $(i,j)\in\alpha$, we denote by $\operatorname{Hook}_{(i,j)}(\alpha)$ the hook centred in position $(i,j)$ and by $\operatorname{h}_{(i,j)}(\alpha)$ the corresponding hook number, i.e. its size.
  • Partitions will be denoted by $\nu=(1^{\nu_{1}},2^{\nu_{2}},\dots)$ and their size is $|\nu|=\sum_{i}i\nu_{i}$.
  • $\Lambda=\bigoplus\Lambda^{d}$ is the graded ring of symmetric functions with coefficients in $\mathbb{Q}$.
@@
  • Power-sum functions are $\mathsf{\mathbf{p}}_{k}(x_{1},x_{2},\dots)=\sum_{i\geq 1}x_{i}^{k}\in\Lambda^{k}$.
@@
Let $\Lambda=\bigoplus\Lambda^{d}$ be the $\mathbb{N}$-graded ring of symmetric functions with coefficients in $\mathbb{Q}$. Two sets of generators of $\Lambda$ are given by Schur functions $\mathsf{\mathbf{s}}_{\alpha}$ and power-sum functions $\mathsf{\mathbf{p}}_{k}$. A Schur function $\mathsf{\mathbf{s}}_{\alpha}$ is homogeneous of degree $|\alpha|$, while a power-sum function $\mathsf{\mathbf{p}}_{k}$ is homogeneous of degree $k$.
Both Schur functions $\mathsf{\mathbf{s}}_{\alpha}$ with $|\alpha|=d$ and monomials in power-sums $\mathsf{\mathbf{p}}^{\mu}$ with $|\mu|=d$ constitute $\mathbb{Q}$-linear bases of each homogeneous component $\Lambda^{d}$. In particular power-sums are algebraically independent generators of
$\Lambda=\mathbb{Q}[\mathsf{\mathbf{p}}_{1},\mathsf{\mathbf{p}}_{2},\dots],$
so that $\Lambda$ may be thought of as a ring of polynomials with rational coefficients in countably many variables $\mathsf{\mathbf{p}}_{1},\mathsf{\mathbf{p}}_{2},\dots$
@@
For any integer $n$, define the derivations
$\mathsf{D}_{n}\colon\Lambda\to\Lambda\,,\qquad\mathsf{D}_{n}=\sum_{k\geq 0}k^{n}\frac{\partial}{\partial\mathsf{\mathbf{p}}_{k}}.$ (2.1)
In [DFO25b] we introduced an analogous derivation $\mathsf{D}_{s}$, where $s$ is a formal variable, and used it to prove recursive formulas that are useful in the computation of tensor products and plethysms of Schur functions. Derivations in (2.1) are obtained by evaluating $\mathsf{D}_{s}$ at $s=n\in\mathbb{N}$.
Remark 2.1.
Notice that the $\mathsf{D}_{n}$ are the derivations $\delta_{\mathsf{q},n}$ we obtain from (1.1), when choosing $q_{n,k}=k^{n}$.
Next, we want to explain how to effectively compute $\mathsf{D}_{n}(\mathsf{\mathbf{s}}_{\alpha})$ using the following well-known combinatorial interpretation, which is explained for instance in [DFO25b, Section 3] (see also [DFO25b, Proposition 2.10]). For each $(i,j)\in\alpha$ we denote by $\alpha_{(i,j)}$ the Young diagram obtained by $\alpha$ by removing the border-strip having the same ends as $\operatorname{Hook}_{(i,j)}(\alpha)$. [^2] Equivalently, we may delete $\operatorname{Hook}_{(i,j)}(\alpha)$ and move the lower component up-left if removing the hook disconnects the diagram.
@@
$\mathsf{D}_{n}(\mathsf{\mathbf{s}}_{\alpha})=\sum_{(i,j)\in\alpha}(-1)^{\varepsilon(i,j)}(\operatorname{h}_{(i,j)})^{n-1}\cdot\mathsf{\mathbf{s}}_{\alpha_{(i,j)}},$
@@
$\mathsf{D}_{n}(\mathsf{\mathbf{s}}_{(3,2,2)})=\mathsf{\mathbf{s}}_{(3,2,1)}+\mathsf{\mathbf{s}}_{2,2,2}+(\mathsf{\mathbf{s}}_{(3,2)}-\mathsf{\mathbf{s}}_{(3,1,1)})2^{n-1}-\mathsf{\mathbf{s}}_{3,1}3^{n-1}+\mathsf{\mathbf{s}}_{1,1,1}4^{n-1}+\mathsf{\mathbf{s}}_{(1,1)}5^{n-1}\,.$
@@
$\mathsf{D}_{n}(\mathsf{\mathbf{s}}_{m})=\sum_{k=1}^{m}k^{n-1}\mathsf{\mathbf{s}}_{m-k}\qquad\mbox{and}\qquad\mathsf{D}_{n}(\mathsf{\mathbf{s}}_{1^{m}})=\sum_{k=1}^{m}(-1)^{k+1}k^{n-1}\mathsf{\mathbf{s}}_{1^{m-k}}.$
@@
If $|\alpha|=n$, then $\mathsf{D}^{\mu}(\mathsf{\mathbf{s}}_{\alpha})=0$ for every $\mu=(1^{\mu_{1}},\dots,\ell^{\mu_{\ell}})$ such that $\mu_{1}+\cdots+\mu_{\ell}>n$. (If we express $\mu$ as a Young diagram, then $\mu_{1}+\cdots+\mu_{\ell}$ is the number of rows of the diagram.)
3. Chern characters of Schur bundles
Throughout this section we assume that $X$ is a connected compact complex manifold and choose the even cohomology ring $\operatorname{H}^{2\ast}(X,\mathbb{Q})$ as the graded commutative algebra $A$ to use in Lemma 1.3. Notice that since $\operatorname{H}^{2n}(X,\mathbb{Q})$ is nonzero for finitely many $n$, then $A$ equals its completion $\widehat{A}$.
3.1. Chern character